What is the magnitude of the relativistic momentum of a proton with a relativistic total energy of 2.9 × 10-10 J?
Answer #1

Relativistic momentum is given by:

P = \gammamv

\gamma = Lorentz factor = 1/sqrt (1 - v^2/c^2)

m = mass of proton = 1.6*10^-19

v = speed of proton

Now given that Total relativistic energy = E = 2.9*10^-10 J

Now we know that E = \gammamc^2

\gamma = E/mc^2 = 2.9*10^-10/(1.67*10^-27*(3*10^8)^2)

\gamma = 1.9295

Now Using above values:

\gamma = 1/sqrt (1 - v^2/c^2)

1 - v^2/c^2 = 1/\gamma^2

v = c*sqrt (1 - 1/\gamma^2)

v = 3*10^8*sqrt (1 - 1/1.9295^2)

v = 2.566*10^8 m/sec

Now Using these values of \gamma and v, relativistic momentum will be:

P = \gamma*m*v

P = 1.9295*1.67*10^-27*2.566*10^8

P = 8.27*10^-19 kg-m/sec

Please Upvote.